OpenStudy (anonymous):

Solve D = 3(2h – k) for h.

OpenStudy (anonymous):

D= 6h-6k (D+6k)/6=h

OpenStudy (texaschic101):

D = 3(2h - k) -- distribute the 3 through the parenthesis D = 6h - 6k -- now add 6k to both sides to get 6h by itself D + 6k = 6h - 6k + 6k -- combine like terms D + 6k = 6h -- now divide by 6 to get k by itself (D + 6k)/6 = 6h/6 (D + 6k)/6 = h